Solve the following, finding all solutions in radians in [0,2pi)
2cos^2(x)-√(3)cos(x)=0

Answer by lwsshak3(11628)

Solve the following, finding all solutions in radians in [0,2)
2cos^2(x)-√(3)cos(x)=0
cosx(2cosx-√3)=0
cosx=0
x=π/2+2πk and 3π/2+2πk, k=integer
and
2cosx-√3=0
cosx=√3/2
x=π/6+2πk and 11π/6+2πk, k=integer (in quadrants I and IV where cos>0)